I have worked this problem several times and just want to be sure I did it correctly. I am supposed to solve with elimination.
2r-5s=-7
5r-2s=55
I came up with the solution of (289/21, 145/21)
Can anyone tell me if this is correct?

Found 2 solutions by Fombitz, scott8148:
Answer by Fombitz(32388)

Yes, those are the correct answers.
Well done!

Answer by scott8148(6628)

multiplying 1st equation by -5 ___ -10r + 25s = 35

multiplying 2nd equation by 2 ___ 10r - 4s = 110

adding resulting equations (eliminating r) ___ 21s = 145 ___ s = 145/21

substituting ___ 2r - 5(145/21) = -7 ___ 42r - 725 = -147 ___ 42r = 578 ___ r = 289/21